# SpatialLLM: A Compound 3D-Informed Design towards Spatially-Intelligent Large Multimodal Models

### v2

@@ Abstract
Humans naturally understand 3D spatial relationships, enabling complex reasoning like predicting collisions of vehicles from different directions. Current large multimodal models (LMMs), however, lack of this capability of 3D spatial reasoning. This limitation stems from the scarcity of 3D training data and the bias in current model designs toward 2D data. In this paper, we systematically study the impact of 3D-informed data, architecture, and training setups, introducing SpatialLLM, a large multi-modal model with advanced 3D spatial reasoning abilities. To address data limitations, we develop two types of 3D-informed training datasets: (1) 3D-informed probing data focused on object’s 3D location and orientation, and (2) 3D-informed conversation data for complex spatial relationships. Notably, we are the first to curate VQA data that incorporate 3D orientation relationships on real images. Furthermore, we systematically integrate these two types of training data with the architectural and training designs of LMMs, providing a roadmap for optimal design aimed at achieving superior 3D reasoning capabilities. Our SpatialLLM advances machines toward highly capable 3D-informed reasoning, surpassing GPT-4o performance by 8.7%. Our systematic empirical design and the resulting findings offer valuable insights for future research in this direction.
@@ 3.3.2 Designing 3D-Informed LMM: A Roadmap
Figure 6: We modernize a standard LLaVA-v1.5 towards the design of a 3D-informed LMM. The bars are the answer accuracies on the SpatialVQA benchmark, indicating the spatial reasoning capability of each model; results for the GPT-4o is shown with the purple bar.
@@ 4.2 Results
Remarkably, our model achieves a performance of 62.7%, outperforming the top proprietary model by 8.7% and the best open-source model by 10.5%. Besides, our model excels in all aspects of spatial reasoning involving 3D distance, 3D orientation and 3D spatial relationship. Moreover, our model not only enhances spatial reasoning performance, but still preserves general visual reasoning (detailed in appendix). Interestingly, although SpatialVLM [14] (implemented in SpaceLLaVA [2]) outperforms other open-source models in overall performance, it falls short in 3D orientation reasoning compared to LLaVA, due to the lack of tuning with 3D orientation-specific data. We exclude SpatialRGPT from this comparison as it relies on additional inputs such as object bounding box coordinates, segmentation masks, and depth maps. In contrast, our SpatialLLM, along with LLaVA and Cambrian, focuses on learning spatial awareness directly from raw images. We will consider models with additional inputs in future work. In sum, our model achieves state-of-the-art performance on spatial reasoning, surpassing API models [1, 4] and also spatial-centric VLM [14].
@@
Upgrading architectures enable better spatial reasoning. In terms of architecture, integrating a mixed vision encoder can improve overall performance especially for the 3D orientation. Then, upgrading the LLM to the more advanced LLama3-8B enhances spatial reasoning performance by 1%, leveraging its improved language reasoning capabilities.
@@
Figure 7: Our model is capable of answering question correctly that needs accurate reasoning on spatial distance and 3D orientation, while GPT-4o either needs more spatial information or gives wrong answer.
Qualitative comparison. Fig. 7 illustrates a scenario involving spatial reasoning. GPT-4o struggles to reason about the spatial relationships between objects due to its limited ability to comprehend their 3D orientation. In contrast, our model successfully understands these spatial relationships and provides accurate answers, demonstrating superior 3D spatial reasoning capability.
@@ C Public Release
Data.
To benefit future research on LMMs with strong 3D spatial reasoning capabilities, we will release all our training and testing data, including the 3D-informed probing data, i.e., 3DI-Pb-IN166K and 3DI-Pb-OI1M, the 3D-informed instruction tuning data, i.e., 3DI-Ft1M, and our testing benchmark, i.e., SpatialVQA.
Code.
All code will be made available.

### v3

@@ Abstract
Humans naturally understand 3D spatial relationships, enabling complex reasoning like predicting collisions of vehicles from different directions. Current large multimodal models (LMMs), however, lack of this capability of 3D spatial reasoning. This limitation stems from the scarcity of 3D training data and the bias in current model designs toward 2D data. In this paper, we systematically study the impact of 3D-informed data, architecture, and training setups, introducing SpatialLLM, a large multi-modal model with advanced 3D spatial reasoning abilities. To address data limitations, we develop two types of 3D-informed training datasets: (1) 3D-informed probing data focused on object’s 3D location and orientation, and (2) 3D-informed conversation data for complex spatial relationships. Notably, we are the first to curate VQA data that incorporate 3D orientation relationships on real images. Furthermore, we systematically integrate these two types of training data with the architectural and training designs of LMMs, providing a roadmap for optimal design aimed at achieving superior 3D reasoning capabilities. Our SpatialLLM advances machines toward highly capable 3D-informed reasoning, surpassing GPT-4o performance by 8.7%. Our systematic empirical design and the resulting findings offer valuable insights for future research in this direction. Our project page is available here.
@@ 3.3.2 Designing 3D-Informed LMM: A Roadmap
Figure 6: We modernize a standard LLaVA-v1.5 towards the design of a 3D-informed LMM. The bars show answer accuracies on SpatialVQA benchmark, indicating spatial reasoning capability of each model; results for the GPT-4o is shown with the purple bar.
@@ 4.2 Results
Remarkably, our model achieves a performance of 62.7%, outperforming the top proprietary model by 8.7% and the best open-source model by 10.5%. Besides, our model excels in all aspects of spatial reasoning involving 3D distance, 3D orientation and 3D spatial relationship. Moreover, our model not only enhances spatial reasoning performance, but still preserves general visual reasoning (detailed in appendix). Interestingly, although SpatialVLM [14] (implemented in SpaceLLaVA [2]) outperforms other open-source models in overall performance, it falls short in 3D orientation reasoning compared to LLaVA, due to lack of tuning with 3D orientation-specific data. We exclude SpatialRGPT from this comparison as it relies on additional inputs such as object bounding boxes, segmentation masks, and depth maps. In contrast, our SpatialLLM, along with LLaVA and Cambrian, focuses on learning spatial awareness directly from raw images. We will consider models with additional inputs in future work. In sum, our model achieves state-of-the-art performance on spatial reasoning, surpassing API models [1, 4] and also spatial-centric VLM [14].
@@
Upgrading architectures enable better spatial reasoning. In terms of architecture, integrating a mixed vision encoder can improve overall performance especially for 3D orientation. Upgrading the LLM to the more advanced LLama3-8B enhances spatial reasoning performance by 1%, leveraging improved language reasoning capabilities.
@@
Figure 7: Our model answers question correctly that needs accurate reasoning on spatial distance and 3D orientation, while GPT-4o needs more spatial information or gives wrong answer.
Qualitative comparison. Fig. 7 illustrates a scenario involving spatial reasoning. GPT-4o struggles to reason about spatial relationships between objects due to limited ability to comprehend 3D orientations. In contrast, our model correctly predicts spatial relationships and provides accurate answers, showing superior 3D spatial reasoning capability.
@@ C Public Release
See “Data & Code” section on our project page.